OpenStudy (anonymous):

the area of the sector that has a central angle of 120 and a radius of 6 inches

OpenStudy (anonymous):

Area = π(6²) x 120/360 = 12π = 37.70

OpenStudy (mrnood):

the area of a sector is given by r^2 * theta however - theta must be in radians It is evident in this case that 120 deg = 1/3 of a circle area of circle = pi r^2